OpenStudy (omarbirjas):

Will medal (sqrt(250x^16))/sqrt(2x) divide and simplify if possible, show your work.

OpenStudy (omarbirjas):

\[\frac{ \sqrt{250x ^{16}}}{ \sqrt{2x} }\]

OpenStudy (welshfella):

write it im indice form 250^1/2 * (x^16)^1/2 ------------------- 2^1/2 * x^1/2

OpenStudy (welshfella):

first can you simplify (x^16)^1/2 recall the rule of indices (x^a)^b = x^ab

OpenStudy (welshfella):

so (x^16)^1/2 = x^ (16 * 1/2) = ?